# Exploring Non-contrastive Self-supervised Representation Learning for Image-based Profiling

### v1

@@ Abstract
Image-based cell profiling aims to create informative representations of cell images. This technique is critical in drug discovery and has greatly advanced with recent improvements in computer vision. Inspired by recent developments in non-contrastive Self-Supervised Learning (SSL), this paper provides an initial exploration into training a generalizable feature extractor for cell images using such methods. However, there are two major challenges: 1) There is a large difference between the distributions of cell images and natural images, causing the view-generation process in existing SSL methods to fail; and 2) Unlike typical scenarios where each representation is based on a single image, cell profiling often involves multiple input images, making it difficult to effectively combine all available information. To overcome these challenges, we propose SSLProfiler, a non-contrastive SSL framework specifically designed for cell profiling. We introduce specialized data augmentation and representation post-processing methods tailored to cell images, which effectively address the issues mentioned above and result in a robust feature extractor. With these improvements, SSLProfiler won the Cell Line Transferability challenge at CVPR 2025.
@@ 1 Introduction
Cell profiling aims to create meaningful representations of cells, which can be utilized for validating compounds in drug discovery and understanding disease mechanisms [8]. Among various cell profiling methods, image-based profiling using microscope images is the most cost-effective approach for generating high-dimensional representations.
Generating representations without human annotations has long been an important goal in computer vision, and recent advances in Self-Supervised Learning (SSL) have brought this closer. Although previous research has made significant progress in image-based profiling using computer vision techniques [6, 26, 5, 34, 33, 37], obtaining a generalizable feature extractor through self-supervised methods remains an open challenge.
Although SSL has been highly successful with natural image processing, directly applying existing SSL methods to cell profiling is not suitable. The main difficulty arises from the distribution gap between fluorescently dyed microscope cell images and natural images. This gap introduces two challenges: 1) State-of-the-art SSL methods aim to learn representations that remain consistent across different views of objects, resulting in informative and generalizable representations. However, these methods depend on carefully designed image augmentations. Applying these augmentations directly to cell images is not appropriate and degrades performance due to dimensional collapse [21]. 2) Unlike natural images, where representations are generated from a 3-channel image, cell images require generating one representation from multiple images with several channels. These channels, which may include fluorescent and brightfield images, contain lower information density compared to natural images. Effectively extracting and merging information from these channels is therefore crucial.
To address the first challenge, we propose augmentations specifically tailored for cell images. Given that the channels are dependent on fluorescent dyes, we introduce a channel-aware color jitter augmentation. Additionally, we simulate imaging noise through microscope noise augmentation. To enhance the robustness of learned representations considering cell morphology and anisotropy, we include elastic transformations and random rotations.
For the second challenge, to make our model adaptable to varying inputs, we use a single-cell image as input to the feature extractor. Then, we design a sequence of feature post-processing steps to merge outputs from multiple images within the same well, producing a final representation. Specifically, we first perform multi-granularity merging by concatenating the individual image representation with averaged dense representations. To merge representations from different sites within a well, we implement interpolation-based merging. To strengthen the causal relationship between representations and perturbation compounds, we apply a cross-plate concentration scheme.
@@ 2.1 Self-supervised Learning
Initially, SSL aimed at solving pretext tasks designed manually by humans [1, 14, 20, 27, 28, 23, 16]. Later, contrastive learning methods [18, 29, 10, 30, 12] significantly outperformed supervised pretraining approaches on image-level tasks, making SSL a mainstream method for model pretraining. Recently, non-contrastive learning, which aims to achieve consistency across different views [17, 18, 7, 10, 11, 3], has attracted increasing interest due to its better ability to generalize. Non-contrastive methods have been successfully applied in various domains, including video representation learning [25, 24] and medical image processing [2]. Nevertheless, applying SSL to dense prediction tasks [39] or long-tailed scenarios [13] remains challenging. More importantly, how to effectively integrate SSL, especially non-contrastive approaches, into image-based cell profiling remains an open research question.
@@ 3.1 Task Definition
3.2 Data Preprocessing
The original dataset consists of 16-bit TIFF images. To reduce disk space usage and accelerate data loading during training, we first convert images to 8-bit format, aligning them with natural image standards, using:
$I(x,y)=\left\lfloor\frac{I(x,y)-\min(I)}{\max(I)-\min(I)}\times 255\right\rfloor.$
Next, we compute the mean and variance for each channel, denoted by $\bm{\mu}$ and $\bm{\sigma}$. During both training and inference, the input images are normalized as follows:
$\bm{x}_{p}=\frac{\bm{x}_{p}-\bm{\mu}}{\bm{\sigma}}.$
3.3 Self-supervised Model Pretraining
3.3.1 Model Framework
We adopt a classical non-contrastive SSL framework based on a Siamese network, which comprises a student model $f_{s}$ and a teacher model $f_{t}$. Consistent with standard self-distillation methods [7, 38, 31], the student model learns by distilling knowledge from the teacher model, while the teacher model is updated using an Exponential Moving Average (EMA) of the student model parameters. A Vision Transformer (ViT) [15] is used as the backbone architecture.
During pretraining, each position $\bm{x}_{p}$ of an image is treated as a single input unit. Unlike natural images, cell images consist of 8 channels rather than three. Due to significant distribution differences between fluorescent and brightfield channels, and because some datasets only include 5 channels, we separately train two models for these channel types. Their features are then combined during testing. Empirical results indicate that this approach performs better than using all 8 channels simultaneously.
For loss computation, we directly employ the DINO v2 framework without modification. It includes three main components: the $\mathcal{L}_{DINO}$ loss for instance-level information distillation [7], the $\mathcal{L}_{iBOT}$ loss for patch-level information reconstruction [38], and the $\mathcal{L}_{KoLeo}$ loss for regularization [32]. The overall loss function is:
$\mathcal{L}_{total}=\mathcal{L}_{DINO}+\lambda_{1}\cdot\mathcal{L}_{iBOT}+\lambda_{2}\cdot\mathcal{L}_{KoLeo},$
where $\lambda_{1}$ and $\lambda_{2}$ are hyperparameters. Since the loss functions remain unchanged during training, further details are omitted due to space limitations. Readers may refer to the original DINO v2 paper [31] for more details.
@@ 3.3.2 Data Augmentation
and subsequently adjusted for contrast relative to its mean $\mu_{c}$:
@@ 4.2 Main Results
  • ViT-Base-14 and 400 epochs means we increased the backbone model size and training duration, achieving better performance.
  • Cross place alignment refers to the post-processing method described in Sec. 3.4.2. While this method significantly improved results in the challenge, it may lead to overfitting and reduced generalization. Therefore, we recommend using it only for this specific challenge.
@@
+ ViT-Base-14 37.0 $\uparrow$ 1.6 OOM
+ 400 epochs training 38.2 $\uparrow$ 1.2 OOM
@@ 6.2 Elastic Deformation
$\tilde{I}_{c}(u,v)=I_{c}\bigl{(}\tilde{\mathbf{x}}(u,v)\bigr{)},\quad c=1,\dots,C.$

### v2

# Self-supervised Representation Learning with Local Aggregation for Image-based Profiling

@@ Abstract
Image-based cell profiling aims to create informative representations of cell images. This technique is critical in drug discovery and has greatly advanced with recent improvements in computer vision. Inspired by recent developments in non-contrastive Self-Supervised Learning (SSL), this paper provides an initial exploration into training a generalizable feature extractor for cell images using such methods. However, there are two major challenges: 1) Unlike typical scenarios where each representation is based on a single image, cell profiling often involves multiple input images, making it difficult to effectively fuse all available information; and 2) There is a large difference between the distributions of cell images and natural images, causing the view-generation process in existing SSL methods to fail. To address these issues, we propose a self-supervised framework with local aggregation to improve cross-site consistency of cell representations. We introduce specialized data augmentation and representation post-processing methods tailored to cell images, which effectively address the issues mentioned above and result in a robust feature extractor. With these improvements, the proposed framework won the Cell Line Transferability challenge at CVPR 2025.
@@ 1 Introduction
Cell profiling aims to learn meaningful representations of cells that support compound validation in drug discovery and the study of disease mechanisms [8]. Among available methods, image-based profiling from microscopy is the most cost-effective way to produce high-dimensional representations.
Learning representations without human annotations has long been an important goal in computer vision, and recent advances in Self-Supervised Learning (SSL) have brought this closer. Previous research has made significant progress in image-based profiling using computer vision techniques [6, 27, 5, 35, 34, 39], yet obtaining a generalizable feature extractor through self-supervised methods remains an open challenge.
Despite SSL’s success on natural images, directly applying standard SSL methods to cell profiling is not suitable. The core issue is the distribution gap between fluorescence microscopy images and natural images. This gap raises two challenges: 1) Natural images typically yield one representation from a single 3-channel input, whereas cell imaging requires one representation from multiple sites in a well and multiple channels, including fluorescent and brightfield images. Leveraging this extra information is crucial. 2) State-of-the-art SSL methods learn view-invariant representations using carefully designed augmentations. Applying these augmentations to cell images is inappropriate and degrades performance due to dimensional collapse [22].
For the first challenge, since cells within a well belong to the same cell line and receive the same treatment, their images should map to similar points in latent space. Guided by this idea, we propose a new SSL framework with an auxiliary local-aggregation branch that improves cross-site consistency. We then apply feature post-processing to merge outputs from multiple sites within a well into a single representation. The aggregations include multi-granularity merging, interpolation-based merging, cross-site merging, and a cross-plate concentration scheme.
To address the second challenge, we design augmentations tailored to cell images. Because channels correspond to fluorescent dyes, we introduce a channel-aware color jitter. We also simulate imaging noise with a microscope noise augmentation. To improve robustness to cell morphology and anisotropy, we include elastic transformations and random rotations.
@@ 2.1 Self-supervised Learning
Initially, SSL aimed at solving pretext tasks designed manually by humans [1, 15, 21, 28, 29, 24, 17]. Later, contrastive learning methods [19, 30, 10, 31, 12] significantly outperformed supervised pretraining approaches on image-level tasks, making SSL a mainstream method for model pretraining. Recently, non-contrastive learning, which aims to achieve consistency across different views [18, 19, 7, 10, 11, 3], has attracted increasing interest due to its better ability to generalize. Non-contrastive methods have been successfully applied in various domains, including video representation learning [26, 25] and medical image processing [2]. Nevertheless, applying SSL to dense prediction tasks [41, 38, 14] or long-tailed scenarios [13] remains challenging. More importantly, how to effectively integrate SSL, especially non-contrastive approaches, into image-based cell profiling remains an open research question.
@@ 3.1 Task Definition
Figure 1: An illustration of the overall framework.
3.2 Self-supervised Learning with Local Aggregation
While previous SSL frameworks conduct the pretext task mainly by augmenting the input image, they do not fully utilize the properties of cell images. In cell painting, cells within a well are from the same cell line and receive the same treatment. Because these images are phenotypically similar and causally consistent, we argue that they should have similar representations in the latent space, which matches the goal of cell profiling.
Motivated by this idea, we propose an SSL framework that uses local aggregation to improve cross-site consistency. An illustration of the framework is shown in Fig. 1.
3.2.1 Learning Framework
We start with a classical non-contrastive SSL framework based on a Siamese network, which comprises a student model $f_{s}$ and a teacher model $f_{t}$. Consistent with standard self-distillation methods [7, 40, 32], the student model learns by distilling knowledge from the teacher model, while the teacher model is updated using an Exponential Moving Average (EMA) of the student model parameters. A Vision Transformer (ViT) [16] is used as the backbone architecture.
During pretraining, each site in a well is treated as a separate input unit. Unlike natural images, cell images may have up to 8 channels rather than three. Because fluorescent and brightfield channels have very different distributions, and some datasets include only 5 channels, we train two models separately for these channel types and combine their features at test time.
Given the image from a site, we first apply random augmentations to produce two views, $\bm{v}_{1}$ and $\bar{\bm{v}}_{1}$, where $\bar{\bm{v}}_{1}$ is additionally masked following [32]. We then introduce an auxiliary branch for local aggregation. We randomly sample an image from another site within the same well and apply random augmentations to obtain the auxiliary view $\bm{v}_{2}$. Details of the augmentations are provided in the next subsection.
For loss computation, we first employ the DINO v2 losses for intra-image alignment. These include three main components: the DINO loss for instance-level distillation [7], the iBOT loss for patch-level reconstruction [40], and the KoLeo loss for regularization [33]. Additionally, we enforce local aggregation by pulling the image-level representations of $\bm{v}_{1}$ and $\bm{v}_{2}$ closer. The overall loss function is:
$\mathcal{L}_{total}$ $=\mathcal{L}_{DINO}(f_{s}(\bar{\bm{v}}_{1}),f_{t}(\bm{v}_{1}))+\mathcal{L}_{DINO}(f_{s}(\bm{v}_{2}),f_{t}(\bm{v}_{1}))$
$+\lambda_{1}\cdot\mathcal{L}_{iBOT}(f_{s}(\bar{\bm{v}}_{1}),f_{t}(\bar{\bm{v}}_{1}))$
$+\lambda_{2}\cdot\mathcal{L}_{KoLeo}(f_{s}(\bm{v}_{1}),f_{s}(\bm{v}_{2})),$
where $\lambda_{1}$ and $\lambda_{2}$ are hyperparameters. Since the loss functions remain unchanged during training, readers can refer to the original DINO v2 paper [32] for more details.
@@ 3.2.2 Data Augmentation
and subsequently adjusted according to its mean $\mu_{c}$:
@@ 4.2 Main Results
  • Local aggregation refers to the local aggregation technique described in Sec 3.
  • ViT-Base-14 means we increased the backbone model size, achieving better performance.
@@
+ Local aggregation 38.6 $\uparrow$ 3.2 -
+ ViT-Base-14 40.0 $\uparrow$ 1.6 OOM
@@ 6.2 Elastic Deformation
$\tilde{I}_{c}(u,v)=I_{c}\bigl(\tilde{\mathbf{x}}(u,v)\bigr),\qquad c=1,\dots,C.$